# On the arithmetic of polynomial ideals

### v1

@@ 4.1. Atoms and sets of lengths in ${\mathcal{M}{\rm on}}(R)$
using (28), we get $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq i\leq a_{\llbracket 3,r\rrbracket}-a_{2}-\beta$, implying that $a_{J}\geq a_{1}+a_{2}+\beta$. Since $\beta=a_{I}$ for some $I\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ and that $r\geq 4$, using Lemma 4.10, (30) and (29), we find a contradiction as in the previous subcase.
Subcase III: $X^{\alpha},Y^{\alpha},X^{i}Y^{j}\in G(\mathfrak{a})$ and $X^{\beta},Y^{\beta},X^{l}Y^{m}\in G(\mathfrak{b})$, i.e.,
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{\alpha}Y^{\beta},X^{\beta}Y^{\alpha},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{i+\beta}Y^{j},X^{i}Y^{j+\beta},X^{l+\alpha}Y^{m},X^{l}Y^{m+\alpha},X^{i+l}Y^{j+m},\dots\rangle,$ (31)
with $i,j,l,m\in\mathbb{N}$ such that $i+j\geq\alpha$, $1\leq i,j<\alpha$, $l+m\geq\beta$ and $1\leq l,m<\beta$.
@@
From (8) we know that $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\in\tilde{\mathfrak{b}}_{r}$ so at least one of the generators in (31) must divide it. From remarks following (8), and in view of subcases I and II, there must exist $i,j,l,m\in\mathbb{N}$ satisfying conditions above such that
$X^{i+l}Y^{j+m}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}.$
Then,
$i+l\leq a_{\llbracket 3,r\rrbracket}-a_{2},$ (32)
$j+m\leq a_{3}-a_{2}.$ (33)
@@
We claim that then, either $X^{i+\beta}Y^{j}\notin\tilde{\mathfrak{b}}_{r}$ or $X^{l}Y^{m+\alpha}\notin\tilde{\mathfrak{b}}_{r}$.
@@
To prove this assumption, let us assume that $X^{i+\beta}Y^{j}\in\tilde{\mathfrak{b}}_{r}$. In view of Lemma 4.1, one of the generators of $\tilde{\mathfrak{b}}_{r}$ in (10) must divide it. The inequality (33) implies that $j<a_{3}-a_{2}$, hence $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{i+\beta}Y^{j}$. Moreover, if $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ is such that $X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{i+\beta}Y^{j}$, then $a_{J}\leq j<a_{3}-a_{2}$ and this can happen only if $a_{J}=a_{1}$ (if $a_{J}=0$, then $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}=a_{\{1\}\cup\llbracket 3,r\rrbracket}=\alpha+\beta>i+\beta$). Thus, $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}=a_{\llbracket 3,r\rrbracket}\leq i+\beta$, i.e., $i\geq a_{\llbracket 3,r\rrbracket}-\beta$ and, from (32), we get $a_{\llbracket 3,r\rrbracket}-\beta+l\leq i+l\leq a_{\llbracket 3,r\rrbracket}-a_{2}$. Now, using the fact observed at the beginning of CASE 2 that $\beta\leq a_{\{1\}\cup\llbracket 4,r\rrbracket}$, we obtain from the previous chain of inequalities, that $l\leq\beta-a_{2}\leq a_{\{1\}\cup\llbracket 4,r\rrbracket}-a_{2}<a_{\llbracket 4,r\rrbracket}$ ($a_{1}-a_{2}<0$). Moreover, we derive from (33) that $m+\alpha<a_{3}-a_{2}+\alpha$.
@@
At this point we are ready to show that $X^{l}Y^{m+\alpha}\notin\tilde{\mathfrak{b}}_{r}$, thus proving our claim. First of all, $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{l}Y^{m+\alpha}$ since $a_{\llbracket 3,r\rrbracket}-a_{2}>a_{\llbracket 4,r\rrbracket}>l$. Moreover, if for some $\bar{J}\subseteq\{1\}\cup\llbracket 3,r\rrbracket$
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{\bar{J}}}Y^{a_{\bar{J}}}\mid X^{l}Y^{m+\alpha},$ (34)
@@
then $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{\bar{J}}\leq l\leq\beta-a_{2}\leq a_{\{1\}\cup\llbracket 3,r\rrbracket}-\alpha-a_{2}$, implying that $a_{\bar{J}}\geq a_{2}+\alpha$. Then, by Lemma 4.10 we have two possibilities: $a_{\bar{J}}=a_{3}-a_{1}+\alpha$ or $a_{\bar{J}}\geq a_{3}+\alpha$. But (34) also implies that $a_{\bar{J}}\leq m+\alpha<a_{3}-a_{2}+\alpha$ and we reach a contradiction in both the cases above.
@@
Let $m,n\in\mathbb{N}^{+}$ such that $m>n$, and let $I\in{\mathcal{M}{\rm on}}(R)$ with $X^{m},Y^{n}\in G(I)$. Assume $I=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus{R}$. Then:
(12) $\operatorname{\mathsf{m-deg}}{(I)}=\alpha+\beta$, with $\alpha=\operatorname{\mathsf{m-deg}}(\mathfrak{a})\geq 1$ and $\beta=\operatorname{\mathsf{m-deg}}(\mathfrak{b})\geq 1$.
(13) there exist $a,b\in\mathbb{N}^{+}$, with $a\geq\alpha$, $b\geq\beta$, and $a+b=m$ such that $X^{a},Y^{\alpha}\in\ G(\mathfrak{a})$ and $X^{b},Y^{\beta}\in\ G(\mathfrak{b})$.
@@
(13) The minimal generators of $\mathfrak{a}$ are necessarily monomials of the form $X^{a}Y^{c}$, with $a+c\geq\alpha$, while those of $\mathfrak{b}$ are monomials of the form $X^{b}Y^{d}$, with $b+d\geq\beta$. We observe that there must be a minimal generator of $\mathfrak{a}$ of the form $Y^{c}$, for some $c\geq\alpha$. Otherwise, $X$ would divide every element of $\mathfrak{a}$ and thus of $\mathfrak{a}\cdot\mathfrak{b}$, but this is impossible since $Y^{n}\in I$. Similarly, there must be a minimal generator of $\mathfrak{a}$ of the form $X^{a}$ for some positive integer $a\geq\alpha$. Arguing in the same way, $\mathfrak{b}$ also has among its minimal generators a monomial of the form $Y^{d}$ and one of the form $X^{b}$, for some $b,d\geq\beta$. Thus, $Y^{c+d}$ is a generator of $I$, while all its other minimal generators are monomials divisible by $X$. Since $Y^{n}\in I$, by Lemma 4.1, $Y^{c+d}$ must divide it, i.e, it must be $c+d\leq n$. But this can clearly happen only if $c=\alpha$ and $d=\beta$. Analogously, $X^{a+b}$ is a generator of $I$ and so it must divide $X^{m}$, whence $a+b\leq m$. But if $a+b<m$, none of the generators of $I$ divides $X^{a+b}$, so the only possibility is that $a+b=m$. ∎
@@
It is worth noting that for an ideal $I\in{\mathcal{M}{\rm on}}(R)$ with $X^{m},Y^{n}\in\ G(I)$ for some $m,n\in\mathbb{N}^{+}$, in order to check whether $I$ is an atom one needs only to compute a finite number of products $\mathfrak{a}\cdot\mathfrak{b}$. Indeed, by Lemma 4.13, if $I=\mathfrak{a}\cdot\mathfrak{b}$, then $X^{a},Y^{\alpha}\in\ G(\mathfrak{a})$, so every monomial $X^{c}Y^{\gamma}\in\ G(\mathfrak{a})$ satisfies $c\leq a$ and $\gamma\leq\alpha$, and analogously a monomial $X^{d}Y^{\delta}\in\ G(\mathfrak{b})$ satisfies $d\leq b$ and $\delta\leq\beta$.
@@
We already know from [21, Prop. 5.10.1] that $I$ is an atom of $I(R)$, hence also of ${\mathcal{M}{\rm on}}(R)$, when $m=n$. Therefore, by symmetry, it is enough to consider the case $m>n$. If $n=1$, equation (1) ensures the atomicity of $I$ over ${\mathcal{M}{\rm on}}(R)$: if $I=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)$, one among $\mathfrak{a}$ and $\mathfrak{b}$ has min-degree zero and must be the whole ring $R$. Let $n\geq 2$ and assume as a contradiction that there exist proper nonzero monomial ideals of $R$, say $\mathfrak{a},\mathfrak{b}$, such that $I=\mathfrak{a}\cdot\mathfrak{b}$. By Lemma 4.13, we have that $X^{a},Y^{\alpha}\in G(\mathfrak{a})$ and $X^{b},Y^{\beta}\in G(\mathfrak{b})$, where $\alpha=\operatorname{\mathsf{m-deg}}(\mathfrak{a})\geq 1$, $\beta=\operatorname{\mathsf{m-deg}}(\mathfrak{b})\geq 1$, $\alpha+\beta=n$, $a\geq\alpha\geq 1$, $b\geq\beta\geq 1$ and $a+b=m$. It follows that $X^{a}Y^{\beta}\in I$, but this is a contradiction since $m>a$ and $n>\beta$, so none of the generators of $I$ divides it. ∎

### v2

@@ 4.1. Atoms and sets of lengths in ${\mathcal{M}{\rm on}}(R)$
then, using (22), we get $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq i\leq a_{\llbracket 3,r\rrbracket}-a_{2}-\beta$, implying $a_{J}\geq a_{1}+a_{2}+\beta$. Since $\beta=a_{I}$ for some $I\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ and $r\geq 4$, using Lemma 4.11, (24) and (23), we find a contradiction as in the previous subcase.
Subcase III: $\mathfrak{a}\supsetneq\langle X^{\alpha},Y^{\alpha}\rangle$, $\mathfrak{b}\supsetneq\langle X^{\beta},Y^{\beta}\rangle$
@@
One of the minimal generators of $\mathfrak{a}\cdot\mathfrak{b}$ must divide $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\in\tilde{\mathfrak{b}}_{r}$. In view of subcases I and II, there must exist $i,j,k,\ell\in\mathbb{N}$ such that $X^{i}Y^{j}\in G(\mathfrak{a})$, $i+j\geq\alpha$, $1\leq i,j<\alpha$, $X^{k}Y^{\ell}\in G(\mathfrak{b})$, $k+\ell\geq\beta$, $1\leq k,\ell<\beta$ and $X^{i+k}Y^{j+\ell}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$. Thus,
$i+k$ $\leq a_{\llbracket 3,r\rrbracket}-a_{2},$ (25)
$j+\ell$ $\leq a_{3}-a_{2}.$ (26)
@@
We claim that at least one of $X^{i+\beta}Y^{j}$ or $X^{k}Y^{\ell+\alpha}$ does not belong to $\tilde{\mathfrak{b}}_{r}$, contradicting $\tilde{\mathfrak{b}}_{r}=\mathfrak{a}\cdot\mathfrak{b}$.
@@
To prove this claim, suppose that $X^{i+\beta}Y^{j}\in\tilde{\mathfrak{b}}_{r}$. One of the minimal generators of $\tilde{\mathfrak{b}}_{r}$ must therefore divide it. Inequality (26) implies $j<a_{3}-a_{2}$, whence $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{i+\beta}Y^{j}$. Moreover, if $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ is such that $X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{i+\beta}Y^{j}$, then $a_{J}\leq j<a_{3}-a_{2}$ and this can happen only if $a_{J}=a_{1}$ (if $a_{J}=0$, then $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}=a_{\{1\}\cup\llbracket 3,r\rrbracket}=\alpha+\beta>i+\beta$). Thus, $a_{\llbracket 3,r\rrbracket}=a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq i+\beta$, i.e., $i\geq a_{\llbracket 3,r\rrbracket}-\beta$ and, from (25), we get $a_{\llbracket 3,r\rrbracket}-\beta+k\leq i+k\leq a_{\llbracket 3,r\rrbracket}-a_{2}$. As we observed at the beginning of CASE 2 that $\beta\leq a_{\{1\}\cup\llbracket 4,r\rrbracket}$, from the previous chain of inequalities we obtain $k\leq\beta-a_{2}\leq a_{\{1\}\cup\llbracket 4,r\rrbracket}-a_{2}<a_{\llbracket 4,r\rrbracket}$. Moreover, we derive from (26) that $\ell+\alpha<a_{3}-a_{2}+\alpha$.
@@
We conclude the proof of the claim by showing that none of the minimal generators of $\tilde{\mathfrak{b}}_{r}$ divides $X^{k}Y^{\ell+\alpha}\notin\tilde{\mathfrak{b}}_{r}$. We first note that $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{k}Y^{\ell+\alpha}$ since $a_{\llbracket 3,r\rrbracket}-a_{2}>a_{\llbracket 4,r\rrbracket}>k$. Moreover, if
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{\bar{J}}}Y^{a_{\bar{J}}}\mid X^{k}Y^{\ell+\alpha}\text{\ for some\ }\bar{J}\subseteq\{1\}\cup\llbracket 3,r\rrbracket,$ (27)
@@
then $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{\bar{J}}\leq k\leq\beta-a_{2}=a_{\{1\}\cup\llbracket 3,r\rrbracket}-\alpha-a_{2}$, implying $a_{\bar{J}}\geq a_{2}+\alpha$. Then, by Lemma 4.11 we have two possibilities: either $a_{\bar{J}}=a_{3}-a_{1}+\alpha$ or $a_{\bar{J}}\geq a_{3}+\alpha$. Since (27) also implies $a_{\bar{J}}\leq\ell+\alpha<a_{3}-a_{2}+\alpha$, both of these possibilities lead to a contradiction.
@@
Let $m,n\in\mathbb{N}^{+}$ such that $m<n$, and let $I\in{\mathcal{M}{\rm on}}(R)$ with $X^{m},Y^{n}\in G(I)$ and $\operatorname{\mathsf{m-deg}}(I)=m$. Assume $I=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus\{R\}$. Then:
(12) $m=\alpha+\beta$, with $\alpha=\operatorname{\mathsf{m-deg}}(\mathfrak{a})\geq 1$ and $\beta=\operatorname{\mathsf{m-deg}}(\mathfrak{b})\geq 1$.
(13) there exist $a,b\in\mathbb{N}^{+}$, with $a\geq\alpha$, $b\geq\beta$, and $a+b=n$ such that $X^{\alpha},Y^{a}\in\ G(\mathfrak{a})$ and $X^{\beta},Y^{b}\in\ G(\mathfrak{b})$.
@@
(13) Since $G(I)=G(\mathfrak{a}\cdot\mathfrak{b})\subseteq\{fg\colon f\in G(\mathfrak{a}),\ g\in G(\mathfrak{b})\}$, necessarily $X^{i},Y^{a}\in G(\mathfrak{a})$ and $X^{j},Y^{b}\in G(\mathfrak{b})$ for some $i\in\llbracket\alpha,m\rrbracket,\ a\in\llbracket\alpha,n\rrbracket,\ j\in\llbracket\beta,m\rrbracket,\ b\in\llbracket\beta,n\rrbracket$ such that $i+j=m$ and $a+b=n$. It follows that $i=\alpha$ and $j=\beta$. ∎
@@
It is worth noting that when $I$ is a monomial ideal such that $X^{m},Y^{n}\in\ G(I)$ for some $m,n\in\mathbb{N}^{+}$, $\operatorname{\mathsf{m-deg}}(I)=\min\{m,n\}$ and $G(I)\subseteq K[X,Y]$, in order to check whether $I$ is an atom one needs only to compute a finite number of products $\mathfrak{a}\cdot\mathfrak{b}$. Indeed, by Lemma 4.14, if $I=\mathfrak{a}\cdot\mathfrak{b}$, then $X^{\alpha},Y^{a}\in\ G(\mathfrak{a})$ for some $\alpha,a\in\mathbb{N}^{+}$, so every monomial $X^{\gamma}Y^{c}\in\ G(\mathfrak{a})$ satisfies $\gamma\leq\alpha$ and $c\leq a$, and analogously there exist $\beta,b\in\mathbb{N}^{+}$ such that every monomial $X^{\delta}Y^{d}\in\ G(\mathfrak{b})$ satisfies $\delta\leq\beta$ and $d\leq b$.
@@
We already know from [21, Prop. 5.10.1] that $I=\langle X^{m},Y^{n}\rangle$ is an atom of ${\mathcal{I}}{(R)}$, hence also of ${\mathcal{M}{\rm on}}{(R)}$, when $m=n$. Therefore, by symmetry, it is enough to consider the case $m<n$. If $m=1$, equation (1) ensures the atomicity of $I$ in ${\mathcal{M}{\rm on}}(R)$: if $I=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)$, one among $\mathfrak{a}$ and $\mathfrak{b}$ has min-degree zero and must therefore be the whole ring $R$. For $m\geq 2$, assume as a contradiction that there exist proper nonzero monomial ideals of $R$, say $\mathfrak{a},\mathfrak{b}$, such that $I=\mathfrak{a}\cdot\mathfrak{b}$. By Lemma 4.14, we have that $X^{\alpha},Y^{a}\in G(\mathfrak{a})$ and $X^{\beta},Y^{b}\in G(\mathfrak{b})$, where $\alpha=\operatorname{\mathsf{m-deg}}(\mathfrak{a})$, $\beta=\operatorname{\mathsf{m-deg}}(\mathfrak{b})$, $\alpha+\beta=m$, $a\geq\alpha\geq 1$, $b\geq\beta\geq 1$ and $a+b=n$. It follows that $X^{\alpha}Y^{b}\in I$, but this is a contradiction since $m>\alpha$ and $n>b$, so neither of the generators of $I$ divides it. ∎